HUBBLE SPACE TELESCOPE DAILY REPORT #1947 PERIOD COVERED: 0000Z (UTC) 08/21/97 - 0000Z (UTC) 08/22/97 Daily Status Report as of 234/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed NIC/2 7331 (Near-IR Properties of the Bulges of Spiral Galaxies) The Near Infrared Camera (Camera Two) was used to make infrared observations of the bulge in spiral galaxy NGC-5985. Near-IR data at HST resolution are essential for correcting for the effect of dust and will enable us to derive robust nuclear properties for the sample. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Three Sets of WF/PC-2 7621 (WF/PC-2 Cycle 7 Supplemental Darks PT1/3) The WF/PC-2 was used to make supplemental dark calibration images. This set of calibrations is intended to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/CCD 7659 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make dark calibration images in order to update the acquisition bad pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed NIC/1 7451 (Gravitational Lensing by Damped Ly-Alpha Absorbers) The Near Infrared Camera (Camera One) was used to make observations of object PG1320+412 in order to study gravitational lensing in damped Lyman-Alpha absorbers. Damped Lyman-Alphas are thought to arise in the disk of the progenitors of present-day spiral galaxies. The observations were completed as scheduled, and no problems were reported. 1.5 Completed Two Sets of STIS/CCD 7601 (CCD Dark and Bias Monitor) The Space Telescope Imaging Spectrograph (CCD) was used to conduct bias and dark monitoring. Also, observations were made to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were completed as planned, and no anomalies were reported. 1.6 Completed NIC/1/2/3 7327 (Elliptical Galaxy Evolution at High Redshift: IR Imaging of 60 Distant Clusters) The Near Infrared Camera (Cameras One, Two and Three) was used to make infrared observations high redshifted elliptical galaxies in galactic cluster CL1322+302. By measuring the structural properties of high redshift cluster elliptical galaxies, it is hoped to resolve the controversy regarding the age and formation history of bright cluster ellipticals. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed STIS/CCD/MA1 7722 (Lamp Flux Measurements) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make lamp flux calibration measurements to determine the optimal configuration for obtaining science-contemporaneous wavecals. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed STIS/CCD/WF/PC-2 6497 (An Ultra-Deep Study of Comets in the Kuiper Belt) The Space Telescope Imaging Spectrograph (CCD) and WF/PC-2 was used to make observations of the Kuiper Belt ultra-deep object KB1. This object, and about 30 Halley-sized objects like it, have been reported to be detected in the Kuiper belt using HST/WF/PC-2 observations taken during Cycle 4. As described in HSTAR 6156, approximately mid-way through this 22-hour proposal, the guide star acquisition scheduled at 233/161302 defaulted to fine lock backup on FGS-1 only when the scan step limit was exceeded. FHST maps before and after the failure indicated good attitude. The six following re-acquisitions were similarly affected. Eight observations were made during the period and may have been impacted. Otherwise, the observations were completed as planned, and no further anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 2 Scheduled Reacquisitions : 13 Successful : 13 As described in 1.8 above, one acquisition defaulted to fine lock on one FGS only. This also affected the following six re-acquisitions. 2.2 FHST Updates: Scheduled : 20 Successful : 20 2.3 Operations Notes: NICMOS Flight Software 4.08 installation was completed at 233/1714Z. This software corrects "pedestal" (elevated bias levels on the first few frames) and "persistence" (effect of charge left over from bright objects) problems caused by powering off detector output amplifiers during "autoflush". "Autoflush" timing patterns are changed to keep the output amplifiers on during "autoflush". On-orbit testing will occur next week. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH